If tanA and tanB are roots of x^2 - 2x-1=0, then sin^2(A+B) is

Grade:11

1 Answers

Arun
3 years ago
Given,
 
x22x1=0
 
x=21(2)±(2)241(1)
 
x=1+2,x=12
 
tanA=1+2,tanB=12
 
A=tan1(1+2),B=tan1(12)
 
A=6721,B=2221
 
Now,
 
sin2(A+B)
 
=sin2(6721+2221)
 
=sin2(90)
 
=1